OpenStudy (anonymous):

I really need help with this: write a general formula to describe the variation: *A varies directly with x^2 ; A=18(pi) when x=3 *F varies inversely with D^2; F=36 when d=6 *M varies directly with the square of D and inversely with the square root of X; M=24 when D=3 and X=16

OpenStudy (dumbcow):

A = kx^2 --> 18pi = 9k -> k = 2pi A = 2pi x^2 F = k/d^2 --> 36=k/36 -> k = 36^2 =1296 F = 1296/d^2 M = kd^2/sqrt(x) --> 24 = 9k/4 -> k = 96/9 = 32/3 M = 32d^2/3sqrt(x)